# Rigidity of the Borell-Brascamp-Lieb Inequality on Weighted Riemannian Manifolds

### 01 Apr 2026 v1

@@ 4 The Brunn-Minkowski Inequality and its Rigidity
The Brunn-Minkowski inequality on Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$, established in [20] is closely related to the Borell-Brascamp-Lieb inequality, whose rigidity can also be considered as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality.
@@
The fact that on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality shows the equivalence between Borell-Brascamp-Lieb inequality and $\mathrm{Ric}_{m,N}\geq Kg$. $\mathrm{Ric}_{m,N}\geq Kg\Rightarrow$ BBL’ is known, and for the inverse implication, we have that the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality, which by [14] implies the $\mathrm{CD}(K,N)$ condition, equivalent to $\mathrm{Ric}_{m,N}\geq Kg$.
@@ 4.1 Rigidity in the null curvature bound
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$
@@
When $A$ is a point $\{x\}$ and $B$ is a ball centred at $x$, rigidity of the Brunn-Minkowski inequality would imply equality in the Bishop-Gromov volume comparison, which further implies that $B$ is a part of a flat cone due to the main result of [7], with a singularity at $\{x\}$ unless the dimension $n$ of $M$ is $N$. In general, equality in the Brunn-Minkowski inequality with $N>n$ does not necessarily imply the existence of a singularity, when $A$ and $B$ are set to be concentric annuli, the Brunn-Minkowski inequality about such $A$ and $B$ would be able to attain equality on a cone with its tip smoothened.
@@ 4.2 Rigidity in the positive and negative curvature bounds
By the scaling property of the $p$-mean, similar to the one used in the Theorem 4.1(ii) and Theorem 4.2 of [2], we have:
@@
and due to the monotonicity of the function $\beta^{t}_{K,N}(r)$ in $r$, $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$ for a.e. $x\in A$. Thus, by the same proof of Theorem 4.2 of [2], we conclude the following result on rigidity.
@@
When $A$ and $B$ are disjoint, we can apply the same argument constructed in [2] with $\frac{s_{K,N}(tx)}{s_{K,N}(x)}$ replaced by $\beta^{t}_{K,N}(x)$. We set $s:=\Theta_{A,B}=\inf_{(x,y)\in A\times B}d(x,y)=d(x,\mathcal{F}_{1}(x))$ for almost every $x\in A$. Fix $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=s$, and $B_{s}(\mathcal{F}_{1}(x))\subset\bigcup_{y\in B}B_{s}(y)$. By construction, $B_{s}(\mathcal{F}_{1}(x))$ is disjoint with $A$. Fix $s_{0}\in(0,s)$ and for every $0<s^{\prime}<s_{0}$, for $z_{s^{\prime}}\in Z_{\frac{s^{\prime}}{2s}}(x,\mathcal{F}_{1}(x))$, $B_{\frac{s^{\prime}}{2}}(z_{s^{\prime}})$ is disjoint with $A$. This implies that $A$ is porous at $x$, that for every $s^{\prime}$ sufficiently small, there is a point $z_{s^{\prime}}$ of distance $s^{\prime}$ to $x$, such that there is an open ball with radius $\frac{s^{\prime}}{2}$ that is outside $A$ centred at which. In this case $m(A)$ cannot be positive, which is a contradiction to the assumption that $m(A)>0$.
@@ 5 Further Problems
    It is possible to generalise Theorem 2 to Finsler manifolds, on which the curvature-dimension condition is described in [18], by considering the equality of the Finsler interpolation inequality presented in the same paper.
@@
    Following the proof of the Borell-Brascamp-Lieb inequality on metric measure spaces in [1], it is also natural to consider the rigidity of this inequality in a synthetic way, on $\mathrm{CD}(K,N)$ or $\mathrm{RCD}(K,N)$ metric measure spaces.
@@
    Considering the Borell-Brascamp-Lieb inequality together with its rigidity and stability on Lorentzian manifolds with timelike Ricci curvature bounds is also a direction of generalisation, under the framework of optimal transport on Lorentzian manifolds constructed in [16] and [17]. Recently, in [5], Cavalletti-Mondino proved the Brunn-Minkowski inequality on Lorentzian manifolds with timelike Ricci curvature bound, and in [9] Farooqui showed the equivalence between timelike Brunn-Minkowski inequality and timelike Ricci curvature bound.

### 20 Apr 2026 v2

@@ 4 The Brunn-Minkowski Inequality and its Rigidity
The Brunn-Minkowski inequality on Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$, established in (Sturm, 2006) is closely related to the Borell-Brascamp-Lieb inequality, whose rigidity can also be considered as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality.
@@
The fact that on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality shows the equivalence between Borell-Brascamp-Lieb inequality and $\mathrm{Ric}_{m,N}\geq Kg$. $\mathrm{Ric}_{m,N}\geq Kg\Rightarrow$ BBL’ is known, and for the inverse implication, we have that the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality, which by (Magnabosco et al., 2024) implies the $\mathrm{CD}(K,N)$ condition, equivalent to $\mathrm{Ric}_{m,N}\geq Kg$.
@@ 4.1 Rigidity in the null curvature bound
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$ (11)
@@
When $A$ is a point $\{x\}$ and $B$ is a ball centred at $x$, rigidity of the Brunn-Minkowski inequality would imply equality in the Bishop-Gromov volume comparison, which further implies that $B$ is a part of a flat cone due to the main result of (De Philippis and Gigli, 2016), with a singularity at $\{x\}$ if $M$ is complete, unless the dimension $n$ of $M$ is $N$. Whilst it may be difficult to construct nontrivial examples of sets that attain equality of the Brunn-Minkowski inequality on complete smooth Riemannian manifolds with dimension less than $N$, it is possible to attain equality in the Brunn-Minkowski inequality on geodesically convex smooth open manifolds.
Example 1.
Let $M$ be the Euclidean half space $\mathbb{R}^{n}_{+}:=\{x=(x^{1},...,x^{n})\in\mathbb{R}^{n}|x^{n}>0\}$ endowed with the measure $m=\|x\|^{N-n}\mathcal{L}^{n}$. Let $A=B_{r}((0,...,1))$ be a ball of radius $r<1$ centred at (0,,1)(0,...,1) and $B=2A=B_{2r}((0,...,2))$. We have $Z_{t}(A,B)=(1+t)A$, and
$m(Z_{t}(A,B))$ $=$ $\int_{Z_{t}(A,B)}dm=\int_{(1+t)A}\|x\|^{N-n}d\mathcal{L}^{n}(x)=\int_{A}(1+t)^{n}\|(1+t)x\|^{N-n}d\mathcal{L}^{n}(x)$
$=$ $(1+t)^{N}m(A)=\Big((1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}\Big)^{N}.$
@@ 4.2 Rigidity in the positive and negative curvature bounds
By the scaling property of the $p$-mean, similar to the one used in the Theorem 4.1(ii) and Theorem 4.2 of (Balogh and Kristály, 2018), we have:
@@
and due to the monotonicity of the function $\beta^{t}_{K,N}(r)$ in $r$, $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$ for a.e. $x\in A$. Thus, by the same proof of Theorem 4.2 of (Balogh and Kristály, 2018), we conclude the following result on rigidity.
@@
When $A$ and $B$ are disjoint, we can apply the same argument constructed in (Balogh and Kristály, 2018) with $\frac{s_{K,N}(tx)}{s_{K,N}(x)}$ replaced by $\beta^{t}_{K,N}(x)$. We set $s:=\Theta_{A,B}=\inf_{(x,y)\in A\times B}d(x,y)=d(x,\mathcal{F}_{1}(x))$ for almost every $x\in A$. Fix $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=s$, and $B_{s}(\mathcal{F}_{1}(x))\subset\bigcup_{y\in B}B_{s}(y)$. By construction, $B_{s}(\mathcal{F}_{1}(x))$ is disjoint with $A$. Fix $s_{0}\in(0,s)$ and for every $0<s^{\prime}<s_{0}$, for $z_{s^{\prime}}\in Z_{\frac{s^{\prime}}{2s}}(x,\mathcal{F}_{1}(x))$, $B_{\frac{s^{\prime}}{2}}(z_{s^{\prime}})$ is disjoint with $A$. This implies that $A$ is porous at $x$, that for every $s^{\prime}$ sufficiently small, there is a point $z_{s^{\prime}}$ of distance $s^{\prime}$ to $x$, such that there is an open ball with radius $\frac{s^{\prime}}{2}$ that is outside $A$ centred at which. In this case $m(A)$ cannot be positive, which is a contradiction to the assumption that $m(A)>0$.
@@ 5 Further Problems
    It is possible to generalise Theorem 2 to Finsler manifolds, on which the curvature-dimension condition is described in (Ohta, 2009), by considering the equality of the Finsler interpolation inequality presented in the same paper.
@@
    Following the proof of the Borell-Brascamp-Lieb inequality on metric measure spaces in (Bacher, 2010), it is also natural to consider the rigidity of this inequality in a synthetic way, on $\mathrm{CD}(K,N)$ or $\mathrm{RCD}(K,N)$ metric measure spaces.
@@
    Considering the Borell-Brascamp-Lieb inequality together with its rigidity and stability on Lorentzian manifolds with timelike Ricci curvature bounds is also a direction of generalisation, under the framework of optimal transport on Lorentzian manifolds constructed in (McCann, 2020) and (Mondino and Suhr, 2023). Recently, in (Cavalletti and Mondino, 2025), Cavalletti-Mondino proved the Brunn-Minkowski inequality on Lorentzian manifolds with timelike Ricci curvature bound, and in (Farooqui, 2025) Farooqui showed the equivalence between timelike Brunn-Minkowski inequality and timelike Ricci curvature bound.